OpenStudy (anonymous):

how many quart sized cans of paint will Janice need to paint if each quart covers 18 ft squared. the height is 6 the base 1 is 10 base 2 is 8 and the slant height is 9.. its a trapezoid..PLEASE HELP

OpenStudy (anonymous):

please help

OpenStudy (goldphenoix):

So you want to find the area of the trapezoid, since you're painting it. To find the area of the trapezoid, you use the formula: \[\large \large \frac{ Base 1 + Base 2 }{ 2 } \times Height = Area\] So we know everything, just plug it in: \[\large \large \frac{ 10 + 8 }{ 2 } \times 9 = 81\] So the area of the trapezoid is 81 ft^2. So if each paint can cover 18 ft^2, then you can divide 81 by 18. This will give you 4.5 can of paint. Or round it up to 5, since you cannot have half a can of paint.

OpenStudy (goldphenoix):

Let me check if I did anything wrong.

OpenStudy (anonymous):

but my geometry teacher gave us just the anwer and it came out to 3..here let me draw it for you

OpenStudy (anonymous):



OpenStudy (anonymous):

the height is 6 slant height 9 the line at the top is 8 the line at the bottom is 10

OpenStudy (goldphenoix):

The height is 6! I thought it was 9! Oh, the slant height is 9, my bad.

OpenStudy (goldphenoix):

So it would be: \[\large \large \frac{ 10 + 8 }{ 2 } \times 6 = 54.\]

OpenStudy (goldphenoix):

54 divided by 18 is 3. I hope this help! Sorry that I mest up with the height, but you understand how I did it, right?

OpenStudy (goldphenoix):

The slant height is uneeded, you only need the height.

OpenStudy (anonymous):

yees awesome you rock :)

OpenStudy (jazzyfa30):

you forgot the slant height and that doesn't come out to be 3 it comes out to be 9

OpenStudy (goldphenoix):

Thanks. :)

OpenStudy (goldphenoix):

@jazzyfa30 the slant height is uneeded.

OpenStudy (anonymous):

no he did it right my teacher gave me the answer and its 3 lol

OpenStudy (jazzyfa30):

oh okay and @GoldPhenoix I need help with problem and good @brittanyneedshelp im glad you got your answer.